OpenStudy (anonymous):

write the equation of the parabola that meets each set of conditions. The vertex is at (-5,1) and the focus is at (2,1).

OpenStudy (anonymous):



OpenStudy (anonymous):

which was does it face?

OpenStudy (anonymous):

to the right? @satellite73

OpenStudy (anonymous):

yes, that means the y term is squared, not the x term

OpenStudy (anonymous):

my concern is how do i find the p value?

OpenStudy (anonymous):

from \((-5,1)\) to \((2,1)\) is how many units?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

that is p

OpenStudy (anonymous):

so would the equation be (y-1)^2 =7(x+5)?

OpenStudy (anonymous):

no but close

OpenStudy (anonymous):

it is not \[(y-k)^2=p(x-h)\] it is \[(y-k)^2=4p(x-h)\]

OpenStudy (anonymous):

oh ok so it would be 28 not 7

OpenStudy (anonymous):

right want to check it?

OpenStudy (anonymous):

idk how to graph the p either

OpenStudy (anonymous):

the p is not in the graph p is just a length here is a nice picture http://www.wolframalpha.com/input/?i=parabola+%28y-1%29^2%3D28%28x%2B5%29

OpenStudy (anonymous):

so how i would i know how much to spread the graph out?

OpenStudy (anonymous):

not sure how to answer that actually

OpenStudy (anonymous):

they are always wider than you think

OpenStudy (anonymous):

oh ok thats fine thank you. i have another problem but ill close this one and post it up

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

it is determined by the p value, you divide the 28 by 2 and get 14 so you go up the foci 14 n down the foci 14. @satellite73